OpenStudy (anonymous):

Which of the following are solutions to the quadratic equation below? Check all that apply. x^2 - 4x + 4 = 3 a. square root 2+3 b. square root 2-3 c. square root 3+2 d. 6 e. negavtive square root 3 +2 f. 2

OpenStudy (anonymous):

First you want to make sure that the other side of the equation is 0. This would become: \[x ^{2} -4x + 1 = 0.\] You can't simply factor this, so you would have to use the quadratic formula.

OpenStudy (anonymous):

\[-b \pm \sqrt{b^2 - 4ac} \over 2a \]

OpenStudy (anonymous):

A quadratic equation is \[ax^2 + bx + c\], so you would plug in the corresponding numbers to the corresponding places.

OpenStudy (anonymous):

\[(-4)^2 \pm \sqrt{(-4)^2 - (-4)(1)(1)} \over 2(1)\]